# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

### v2

@@ 2 A novel framework for ranking problem based on pairwise comparisons using MWFAS
Based on our analysis of the existing algorithms, we chose to implement the heuristic algorithm from [10] for our ranking algorithm due to its efficiency in practice. The local ratio-based heuristic provides an effective trade-off between solution quality and runtime performance, particularly for graphs with a large number of vertices and edges.
2.1 Highlights of Our Work
  • Speed: Our method is very fast. As shown in the results, our method provides an answer in seconds. [20] provided many tables; however, none of them showed the running time of the experiment. It was only claimed that if an experiment did not answer after one week, its performance in the table was filled with ”NaN”. [15] and [26] also did not provide the running time of their programs. To the best of our knowledge, our method is the fastest in this regard, and its performance is comparable to that of other methods.
  • Flexibility: Our method uses an algorithm to solve the MWFAS problem as a black box. The user can choose any algorithm for solving MWFAS, thus allowing them to handle the trade-off between running time and performance. If performance is more important, more precise algorithms can be used; otherwise, faster heuristics can be chosen.
  • Simplicity: Our method is very simple, using only well-known algorithms like topological sort, depth-first search (or breadth-first search) to find cycles, and heap data structures.
@@ 2.2 Algorithm
Input: Pairwise comparisons of elements as a weighted directed graph file.
Output: Rankings of elements.
Step 1: Read Graph and Initialize Structures;
Parse the input file to read edges $\{(u,v,w)\}$ representing directed edges with weights $w$;
Initialize an adjacency list representation for the graph;
Store edge weights in a dictionary;
Step 2: Find Minimum Weighted Feedback Arc Set (MWFAS);
while there is a cycle in the graph do
       Identify a cycle using depth-first search;
       Compute the minimum weight among edges in the cycle;
@@
             Reduce its weight by the minimum weight;
             if its weight becomes zero then
                   Remove the edge from the graph;
Sort removed edges based on their weight in decreasing order;
foreach edge $(u,v)$ removed in Phase 1 do
       if re-adding $(u,v)$ does not create a cycle then
             Re-add $(u,v)$ to the graph;
@@
Perform a topological sort of the modified graph;
Assign rankings to vertices based on their positions in the topological order;
Break ties using (the sum of outgoing edge weights- the sum of incoming edges weights)/(total number of incoming and outgoing edges) ;
return Final rankings and metrics (e.g., removed edges and weights);
Algorithm 1 Rank Elements Based on Pairwise Comparisons
Algorithm 1 is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach. The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management. The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG). We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles. The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights. By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
@@ 3 Experimental Results
The results from our algorithm simulation provide an insightful comparison to the methods discussed in [20]. Below, we analyze these results in detail:
Efficiency and Speed: Our results, as presented in the ”Time (Seconds)” column, demonstrate the efficiency of our approach across all datasets. With an average computation time between 0.02 and 0.35 seconds, our algorithm shows significant advantages in terms of speed. Compared to the methods mentioned in [20], such as [2] and [16], which can be computationally expensive, our algorithm provides substantially faster performance even without leveraging deep learning frameworks.
@@
Loss Metrics: The ”Naive Loss” and ”Simple Loss” values indicate the quality of the rankings produced by our algorithm. In [20], differentiable loss functions such as ”upset, naive” and ”upset, ratio” are introduced to minimize ranking upsets. When compared to these metrics, our ”Naive Loss” and ”Simple Loss” values are both competitive and consistently low. For example, the average ”Simple Loss” across datasets often remains below 1.0, suggesting that our heuristic-based approach maintains strong agreement with the pairwise comparisons, while also being much simpler to implement and computationally efficient.
This is the exact definition of naive upset loss, simple upset loss, and ratio upset loss based on [20]:
  • Upset Naive: The naive upset loss, denoted as $L_{upset,naive}$, is defined to measure the fraction of pairwise relationships whose inferred ranking is in the opposite direction to what the original pairwise measurement indicates. Formally, given the definitions:
@@
    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is a skew-symmetric matrix representing the difference between the pairwise comparisons.
    • $T_{1}=r1^{T}-1r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ represents the ranking scores of the items, and $T_{1}$ represents the difference in ranking scores for each pair of items.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The naive upset loss is given by:
@@
    $L_{upset,naive}=\sum_{i,j:M_{1}(i,j)\neq 0}\left(\text{sign}(T_{1}(i,j))\neq\text{sign}(M_{1}(i,j))\right)/t.$ (1)
@@
  • Upset Simple: The simple upset loss, denoted as $L_{upset,simple}$, extends the naive upset loss by differentiating between ties and opposite signs. Given the sign function acting element-wise on the corresponding entries, the simple upset loss is defined as:
    $L_{upset,simple}=\frac{\lVert\text{sign}(T_{1})-\text{sign}(M_{1})\rVert^{2}_{F}}{t},$ (2)
    where $t$ represents the number of nonzero elements in $M_{1}$. Unlike the naive upset, this metric distinguishes between ties and cases where the predicted signs are opposite.
  • Upset Ratio: The ratio upset loss, denoted as $L_{upset,ratio}$, takes into account not only the sign but also the magnitude of the rankings. It normalizes the differences in ranking scores and pairwise comparison weights to make them comparable. The ratio upset loss is given by:
    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2},$ (3)
    where:
    • $T_{2}=r1^{T}+1r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise score differences, and $\epsilon>0$ is a small positive value used for numerical stability.
    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise comparison weights.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The ratio upset loss considers both the direction and magnitude of the relationships between nodes and helps to account for more nuanced comparisons that are not purely binary in nature.
Input: $adjacency\_matrix$, $scores$, $index$, $lower\_bound$, $upper\_bound$, $epsilon$, $steps$
Output: Optimal score value for $index$
for $step\leftarrow 1$ to $steps$ do
       $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3}$;
       $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3}$;
@@
      if $upper\_bound-lower\_bound<epsilon$ then
@@
$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2}$;
@@
Algorithm 2 Trinary Search Optimization for Minimizing Ratio Upset Loss
Input: $adjacency\_matrix$, $scores$,$num\_iterations$
Output: Optimized $scores$
for $iteration\leftarrow 1$ to $num\_iterations$ do
@@
       for $i\leftarrow 1$ to length$(sorted\_indices)$ do
@@
            $scores[index]\leftarrow\textsc{TrinarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
              /* Trinary search optimization */
return $scores$;
@@
3.1 Trying to decrease ratio loss without increasing naive loss or simple loss
In experiments, we observed that while our algorithm’s naive loss and simple loss are better than the state-of-the-art, it is not the case for the ratio loss. Therefore, we designed an algorithm to decrease the ratio loss without changing the ranking order. We should note that the ratio loss depends not only on the ranking of the elements but also on the scores we assign to the elements. Therefore, we designed Algorithm 3 to update the scores of the elements without changing the rankings. Without loss of generality, we can consider that $scores_{1}\leq scores_{2}\leq...\leq scores_{n}$. In each algorithm step, it iterates over all of the scores. For element $i$, the algorithm does a trinary search between $scores_{i-1}$ and $scores_{i+1}$ to find the optimal amount for $scores_{i}$. Please note that because still $scores_{i}$ will remain between $scores_{i-1}$ and $scores_{i+1}$, naive loss and simple loss do not change, but ratio loss may be decreased. The reason that trinary search works well is that based on our experiments, with a very good approximation, we can consider that ratio loss function for $scores_{i}$ from $scores_{i-1}$ to $scores_{i+1}$, when other elements of $scores$ array do not change, is a unimodal function. The results of using  Algorithm 2 for 40 iterations can be seen for some datasets in  Table 5 It is crucial to note that maybe no solution exists that is better in all three loss functions than the GNNRank algorithm.
@@ 3.1 Trying to decrease ratio loss without increasing naive loss or simple loss
3.2 Analysing the results
Robustness Across Different Datasets: Our approach shows robustness across various types of datasets, including sports (England Football datasets) and academic hiring (Business FM and Computer Science datasets). While [20] emphasizes leveraging trainable similarity matrices to adapt to different graph types, our method, while simpler, consistently achieves low losses and high computational efficiency. This suggests that our heuristic is versatile and performs well irrespective of the dataset characteristics, which aligns with the goal of creating a flexible ranking framework, as described in [20].
Comparative Insights: [20] introduces inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, especially on more complex datasets. Our approach, although not relying on such advanced machine learning techniques, delivers comparable ranking quality while significantly outperforming these methods in terms of computation time. This highlights an important trade-off between using graph neural networks for adaptive learning versus employing efficient heuristic-based methods that achieve similar or better results in specific cases.
Lets have a look to the execution times for [20] commands. We run some commands that are written on the formal Github repository of [20] in link https://github.com/SherylHYX/GNNRank/blob/main/README.md. Command python ./train.py --all_methods DIGRAC syncRank --dataset animal creates a GNNRank model for animal data using DIGRAC as GNN, also produces results on SyncRank. Executing it took 47 seconds. Command python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 creates a GNNRank model for ERO data using both DIGRAC and ib as GNN with 350 nodes, using 0.05 as learning rate. Each epoch in GNNRank has two splits. After 90 seconds, this command had done one split from each of three epochs. The default number of epochs in GNNRank is 1000 epochs. Command python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter creates a GNNRank model for basketball data in season 2010 using all baselines excluding mvr and also saves predicted results. After 90 seconds, it had done one split from each of seven epochs.
In conclusion, our results suggest that our approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by GNN-based methods. Our method’s loss metrics align well with those reported in [20], while providing a notable advantage in processing speed, making it ideal for real-time or large-scale ranking applications.
@@ 3.3 Designing Appropriate loss functions
We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works. Probably better loss functions can be defined for measuring rankings without needing a truth ground. One of them is the sum of the weight of the violated edges(comparisons) divided by some of the weight of all edges. Its computation can be done in $O(E)$ when $E$ is the number of edges of the graph.
Another Possible loss function is $\frac{\sum_{\forall(i,j)|rank(i)>rank(j)}(w_{ij}*(rank(i)-rank(j))}{\sum_{\forall(i,j)}(w_{ij}*(|rank(i)-rank(j)|)}$. This formula gives a coefficient to each edge. The coefficient shows how the ranking uphold(violate) that edge. The more distance the two endpoints of an edge have from each other in the ranking, the more that edge is upholded(violated).

### v3

@@ 2 A Framework for Ranking from Pairwise Comparisons via MWFAS
Based on our review of existing algorithms, we adopt the heuristic proposed by Chakrabarti and Imamura [10] as the core method for our ranking framework, owing to its practical efficiency. The local-ratio based heuristic provides a favorable balance between solution quality and runtime, particularly for graphs with a large number of vertices and edges.
2.1 Highlights of Our Approach
  • Speed: Our method is highly efficient in practice, typically producing results within seconds on the benchmark datasets used in our experiments. In contrast, He et al. [20] reported extensive performance tables but did not include runtime measurements, noting only that experiments exceeding one week were marked as “NaN.” Other recent works, such as [15] and [26], also did not report computational runtime. To the best of our knowledge, our approach is among the fastest available for this task, while achieving competitive ranking performance.
  • Flexibility: Our framework treats the MWFAS solver as a modular component, allowing users to select any algorithmic strategy to handle the feedback arc set problem. This design enables flexible control over the trade-off between computational efficiency and solution accuracy: more precise algorithms may be employed when quality is paramount, whereas faster heuristics can be used when scalability is the primary concern.
  • Simplicity: Our method is conceptually simple and relies on well-established graph algorithms, including topological sorting and cycle detection via depth-first (or breadth-first) search, along with standard heap-based data structures. This simplicity contributes to ease of implementation and facilitates extensibility.
@@ 2.2 Algorithm
Input: A weighted directed graph representing pairwise comparisons.
Output: A ranking of the vertices.
Step 1: Read Graph and Initialize Data Structures;
Parse the input file to construct the set of weighted directed edges $\{(u,v,w)\}$;
Initialize an adjacency list representation of the graph;
Store edge weights in a suitable data structure;
Step 2: Compute a Minimum Weighted Feedback Arc Set (MWFAS);
while the graph contains a directed cycle do
   Identify a directed cycle using depth-first search;
   Let $\Delta$ be the minimum weight among edges in the cycle;
@@
      Decrease $w_{uv}$ by $\Delta$;
    if $w_{uv}=0$ then
         Remove $(u,v)$ from the graph;
Store all removed edges;
Sort removed edges in decreasing order of their weights;
foreach edge $(u,v)$ in the sorted list do
 if adding $(u,v)$ back does not create a directed cycle then
      Reinsert $(u,v)$ into the graph;
@@
Perform a topological sort of the resulting directed acyclic graph;
Assign rankings according to the topological order;
Break ties using the score
$\frac{\sum_{(i,j)\in E}w_{ij}-\sum_{(j,i)\in E}w_{ji}}{\deg^{+}(i)+\deg^{-}(i)},$
where $\deg^{+}(i)$ and $\deg^{-}(i)$ denote the out-degree and in-degree of $i$, respectively;
return the final ranking and relevant metrics (e.g., removed edges and their weights);
Algorithm 1 Ranking from Pairwise Comparisons via MWFAS
Algorithm 1 implements a straightforward procedure for solving the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) framework. The input graph is represented using an adjacency list along with a weight mapping, enabling efficient traversal and edge-weight updates. The algorithm begins by reading the input graph and iteratively removing cycles via depth-first search (DFS), ensuring that the resulting structure becomes a directed acyclic graph (DAG). DFS is also used to evaluate whether removed edges can be reinserted without reintroducing cycles, thereby maintaining connectivity whenever possible. A heap-based topological sort is employed to generate a ranking of the vertices, with ties resolved using a weight-based scoring function. By combining DFS with heap data structures, the algorithm remains both conceptually simple and computationally efficient, offering a practical approach for handling large graphs while minimizing overhead.
@@ 3 Experimental Results
The results obtained from our algorithm provide a meaningful comparison to the methods evaluated in [20]. Below, we summarize and analyze the key findings.
Efficiency and Speed: The computation times reported in the “Time (Seconds)” column demonstrate the efficiency of our approach across all datasets. With average runtimes between $0.02$ and $0.35$ seconds, the algorithm exhibits substantial advantages in terms of speed. In contrast, methods referenced in [20], including [2] and [16], are known to be computationally expensive, and the authors do not report explicit runtime measurements. Notably, [20] indicates that experiments exceeding one week were terminated and marked as “NaN,” suggesting significant scalability limitations. Our results show that competitive performance can be achieved using a lightweight, learning-free heuristic without reliance on deep learning frameworks.
@@
Loss Metrics: The “Naive Loss” and “Simple Loss” values characterize the quality of the rankings generated by our algorithm. He et al. [20] introduced differentiable loss functions—including “upset, naive,” “upset, ratio,” and “upset, margin”—to quantify ranking inconsistencies and optimize rank assignments. In comparison, our loss values are consistently low across datasets. For instance, the average “Simple Loss” frequently remains below $1.0$, indicating strong agreement with the pairwise comparisons, despite the use of a heuristic rather than a differentiable approach. These results suggest that a combinatorial method can achieve competitive ranking quality while being substantially easier to implement and more computationally efficient.
To facilitate comparison with [20], we reproduce the formal definitions of the relevant loss functions below:
  • Upset Naive: The naive upset loss, denoted $L_{\text{upset, naive}}$, measures the fraction of pairwise relationships whose inferred ranking contradicts the observed pairwise comparison. Let
@@
    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is its induced skew-symmetric form,
    • $T_{1}=r\mathbf{1}^{T}-\mathbf{1}r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ is the vector of ranking scores,
    • $t$ denote the number of nonzero elements in $M_{1}$.
    Then the naive upset loss is defined as
@@
    $L_{\text{upset, naive}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\mathbb{I}\!\left[\operatorname{sign}(T_{1}(i,j))\neq\operatorname{sign}(M_{1}(i,j))\right],$ (1)
    where $\mathbb{I}(\cdot)$ is the indicator function.
@@
  • Upset Simple: The simple upset loss, denoted $L_{\text{upset, simple}}$, extends the naive loss by distinguishing between ties and opposite signs. It is defined as
    $L_{\text{upset, simple}}=\frac{\|\,\operatorname{sign}(T_{1})-\operatorname{sign}(M_{1})\,\|^{2}_{F}}{t},$ (2)
    where $\|\cdot\|_{F}$ denotes the Frobenius norm. Unlike the naive loss, this metric penalizes both disagreement in direction and differences involving ties.
  • Upset Ratio: The ratio upset loss, denoted $L_{\text{upset, ratio}}$, incorporates both sign and magnitude information by normalizing pairwise score differences and comparison weights. Let
    • $T_{2}=r\mathbf{1}^{T}+\mathbf{1}r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ be a normalization matrix for score differences,
    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ be a normalization matrix for comparison weights,
    • $t$ denote the number of nonzero elements in $M_{1}$,
    • $\epsilon>0$ a small constant for numerical stability.
    Then the ratio upset loss is defined as
    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2}.$ (3)
    This metric captures both directional disagreement and magnitude mismatch, enabling more nuanced evaluation of pairwise relationships than purely binary losses.
Input: adjacency_matrixadjacency\_matrix, scoresscores, indexindex, lower_boundlower\_bound, upper_boundupper\_bound, epsilonepsilon, stepssteps
Output: Optimal score value for indexindex
for $step\leftarrow 1$ to stepssteps do
 $mid1\leftarrow lower\_bound+\frac{upper\_bound-lower\_bound}{3.0}$;
 $mid2\leftarrow upper\_bound-\frac{upper\_bound-lower\_bound}{3.0}$;
@@
 else if $loss1>loss2$ then
    $lower\_bound\leftarrow mid1$;
@@
    $upper\_bound\leftarrow mid2$;
 if upper_boundlower_bound<epsilonupper\_bound-lower\_bound<epsilon then
@@
$scores[index]\leftarrow\frac{lower\_bound+upper\_bound}{2.0}$;
@@
Algorithm 2 Ternary Search Optimization for Minimizing Ratio Upset Loss
Input: adjacency_matrixadjacency\_matrix, scoresscores,num_iterationsnum\_iterations
Output: Optimized scoresscores
for $iteration\leftarrow 1$ to num_iterationsnum\_iterations do
@@
 for $i\leftarrow 1$ to length(sorted_indices)(sorted\_indices) do
@@
    $scores[index]\leftarrow\textsc{TernarySearchOptimize}(adjacency\_matrix,scores,index,lower\_bound,upper\_bound,\epsilon,\text{steps})$;
    /* Ternary search optimization */
return scoresscores;
@@
3.1 Reducing Ratio Loss Without Increasing Naive or Simple Loss
In our experiments, we observed that although our algorithm achieves lower naive and simple loss than state-of-the-art methods, its ratio loss can be relatively higher. Motivated by this observation, we designed a post-processing procedure to reduce ratio loss without altering the ranking order. It is important to note that the ratio loss depends not only on the relative ordering of elements but also on the numerical scores assigned to them. Consequently, we developed Algorithm 3, which updates the scores while preserving their ordering.
Without loss of generality, let the scores satisfy $scores_{1}\leq scores_{2}\leq\cdots\leq scores_{n}$. In each iteration, the algorithm sweeps over the score vector and, for each element $i$, performs a ternary search over the interval $scores_{i-1},scores_{i+1}]$ to determine an updated value for $scores_{i}$. Since the updated value remains between its two neighbors, the induced ranking does not change, and therefore the naive and simple loss metrics remain unchanged, while the ratio loss may be decreased.
Empirically, we found that the ratio loss as a function of $scores_{i}$ (with all other scores fixed) is well-approximated by a unimodal function over the interval $scores_{i-1},scores_{i+1}]$. This observation justifies the use of ternary search, which provides an efficient strategy for identifying a near-optimal value for $scores_{i}$. The performance of Algorithm 3 using Algorithm 2 for 40 iterations is reported for selected datasets in Table 5.
It is also important to note that there may not exist a scoring vector that simultaneously outperforms GNNRank on all three loss metrics. Therefore, improvements in ratio loss should be interpreted as complementary rather than universally dominant.
@@ 3.1 Reducing Ratio Loss Without Increasing Naive or Simple Loss
3.2 Analyzing the Results
Robustness Across Datasets: Our approach demonstrates robustness across diverse datasets, including sports datasets (e.g., England football) and academic hiring datasets (e.g., Business FM and Computer Science). While He et al. [20] emphasize learning trainable similarity matrices to adapt to different graph structures, our simpler heuristic consistently achieves low loss values and high computational efficiency. This suggests that the method performs reliably under varying data characteristics, aligning with the goal of designing a versatile ranking framework, as highlighted in [20].
Comparative Insights: He et al. [20] incorporate inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, particularly on complex datasets. In contrast, our approach does not rely on such machine-learning-based components, yet delivers comparable ranking quality while substantially reducing computation time. This illustrates an important trade-off between adaptive learning mechanisms and efficient heuristic methods, where the latter can yield competitive results with significantly lower computational overhead.
Runtime Observations: To provide a clearer comparison of computational cost, we executed several commands from the official implementation of [20], publicly available at https://github.com/SherylHYX/GNNRank. Representative examples include:
  • python ./train.py --all_methods DIGRAC syncRank --dataset animal trains a GNNRank model on the animal dataset using DIGRAC and the SyncRank baseline. Execution required approximately $47$ seconds.
  • python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 trains a GNNRank model on a synthetic ERO dataset with $350$ nodes using DIGRAC and ib. After approximately $90$ seconds, the process completed one split from each of three epochs. The default configuration for GNNRank uses $1000$ epochs.
  • python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter trains on the basketball (2010) dataset using several baselines (excluding mvr). After roughly $90$ seconds, one split from each of seven epochs had been completed.
These results indicate that training GNNRank models generally incurs substantial computational cost, particularly given default settings of hundreds or thousands of epochs. In contrast, our heuristic produces rankings in under a second on all datasets considered, without requiring model training or hyperparameter tuning.
Summary: Overall, our results suggest that the proposed approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by graph neural network–based methods. While our loss metrics are generally aligned with those reported in [20], our substantially lower runtime makes the method attractive for real-time or large-scale ranking applications.
@@ 3.3 Designing Appropriate Loss Functions
The naive and simple upset losses used in [20] provide convenient benchmarks for comparison with prior work. However, neither metric is well suited for evaluating rankings based on weighted pairwise comparisons, as they treat all upsets equally and do not account for the magnitude of disagreement. We report these metrics primarily for consistency with existing literature, but alternative loss functions may offer a more principled assessment of ranking performance when edge weights represent meaningful strengths.
One natural alternative is to measure the total weight of violated edges normalized by the total weight of all comparisons:
$L_{\text{weighted}}=\frac{\sum_{(i,j):\,\pi(i)>\pi(j)}w_{ij}}{\sum_{(i,j)}w_{ij}},$
where $\pi(\cdot)$ denotes the ranking function. This metric directly quantifies the proportion of weighted information contradicted by the ranking and can be computed in $O(|E|)$ time.
A second alternative incorporates both whether an edge is upheld and the distance between the endpoints in the ranking:
$L_{\text{margin}}=\frac{\sum_{(i,j):\,\pi(i)>\pi(j)}w_{ij}\,(\pi(i)-\pi(j))}{\sum_{(i,j)}w_{ij}\,|\pi(i)-\pi(j)|}.$
Here, each edge contributes proportionally to its importance ($w_{ij}$) and the degree to which the ranking respects or violates it. Larger separations between correctly ordered vertices reduce the numerator, while large separations between misordered vertices are penalized more heavily. Thus, this loss function rewards rankings that not only order vertices correctly but also separate them meaningfully.
Designing loss functions that explicitly reflect ranking quality under weighted comparisons—without requiring access to ground truth—remains an important topic for future work.